OpenStudy (rsadhvika):

what are the limit points of below set in \(\mathbb{R}\) \[A = \left\{\frac{1}{n} : n\in \mathbb{N}\right\} \]

OpenStudy (fibonaccichick666):

Well, what is the biggest that function can be?

OpenStudy (rsadhvika):

i think it is bounded by (0, 1)

OpenStudy (rsadhvika):

** (0, 1]

OpenStudy (fibonaccichick666):

mhmm, I would maybe use inf and sup, if it is a set theory class

OpenStudy (fibonaccichick666):

infimum, and supremum-please forgive the spelling

OpenStudy (rsadhvika):

sup is 1 inf is 0

OpenStudy (fibonaccichick666):

yep, and does it attain it's sup? inf?

OpenStudy (rsadhvika):

it attains sup when n=1

OpenStudy (fibonaccichick666):

good, and inf?

OpenStudy (rsadhvika):

nope 0 side is open

OpenStudy (fibonaccichick666):

yep, then you got it

OpenStudy (fibonaccichick666):

I would just be very specific when you write it down

OpenStudy (rsadhvika):

\[\large \left\{\frac{1}{1},\frac{1}2{}, \cdots \right\}\]

OpenStudy (rsadhvika):

okay thanks!

OpenStudy (fibonaccichick666):

np, and if you have (as it appears) used the terms inf and sup, definitely use those to describe. Also, make sure you state which attains :) have a good day!